OpenStudy (anonymous):

A train traveling 50 mph left a station 30 minutes before a second train running at 55 mph. How soon did the second train overtake the first? If 50x represents the distance the slower train travels, then the faster train travels is it 55(x + 0.5)?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

The faster train's travel time is half an hour less than the slower train's. The "plus" sign should be a "minus" sign.

OpenStudy (anonymous):

if s is the distance then s = 50t and for second train s = 55(t - 0.5) thus 50t = 55t - 27.5 5t = 27.5 t = 5.5 hours

OpenStudy (anonymous):

The distance the faster train is 55(x - .5).

OpenStudy (anonymous):

yep